A body of mass 5kg is acted upon by two perpendicular forces of magnitude 8N and 6N. Find the magnitude and direction of the acceleration

Find magnitude and direction
Answer #1

Net force, F 2=82+62  N

              F2=100 N   and hence, F= 10N

Acceleration, a = F/m= 10/5= 2m/s2

If the net force F makes an angle beta(b)  with the horizontal, then the acceleration will also make an angle beta (b) with the horizontal. this is because the acceleration is always in the direction of the force.

Tan b = 6/8 =3/4 = 0.75

Hence, b = tan-1(0.75)= 36.9o

 

answered by: pslaw